Use the formula for the general term (the nth term) of a geometric sequence to find the indicated term of the sequence with the given first term, a1, and common ratio, r.
Find a6 when a1 = 7, r = 2.

Answer by stanbon(75887)

Use the formula for the general term (the nth term) of a geometric sequence to find the indicated term of the sequence with the given first term, a1, and common ratio, r.
Find a6 when a1 = 7, r = 2.
---
A(6) = a(1)*r^5 = 7*2^5 = 7*32 = 224
---------------------------
Cheers,
Stan H.